If the national debt of a country (in trillions of dollars)tyears from now is given by the indicated function, find the relative rate of change of the debt12years from now. (Round your answer to two decimal places.)

N(t) =0.6+1.3e0.01t

1
Expert's answer
2022-03-29T12:11:53-0400

"N(t) =0.6+1.3e^{0.01t}\\ N(12)=0.6+1.3e^{0.01\times12}\\N(12)=2.07t"